A2 + B2 = C2
A=6'
B=8'
C=10'
How did they arrive at c = 10'

Answer by arunpaul(104)

A^2 + B^2 = C^2
^ indicate square if the variable
given
A = 6'
B= 8'
putting the value of A and B in the given equation we have
or 6^2 + 8^2 = C^2
or 36 + 64 = C^2
or 100 = C^2
or 10 = C
or C = 10 (Ans)